How much force does an 95.0kg astronaut exert on his chair while sitting at rest on the launch pad?


How much force does the astronaut exert on his chair while accelerating straight up at 11.0m/s2 ?
Answer #1

a) 95 *9.81 = 931.95 N


b) 95 * (9.81 + 11 ) = 1976.95N
Answer #3

at rest only gravity acts

so 95*9.8=931N

upward acceleration=11m/s2

net acceleration=(11+9.8)=20.8m/s2

Force=95*20.8=1976N